# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 1.3 Characterization of Traces
Because the value of $T$ must be equal on all conjugacy classes of $\mathbb{D}_{n}$, we have
$T(e_{q}h)$ $=\frac{1}{n}T\left(\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{i}h\right)$
$=\frac{1}{n}\left(\sum\limits_{i=0}^{m-1}\varepsilon^{2iq}\right)T(h)+\frac{1}{n}\left(\sum\limits_{i=0}^{m-1}\varepsilon^{(2i+1)q}\right)T(gh)$
@@
The following theorem classifies traces on $H_{c}$. In particular, Conditions (0)(0) express trace values on $H_{c}$ in terms of trace values on elements of the form $R(k)e_{q}$, with $R$ a polynomial, and Conditions (0) and (0) restrict possible trace values on such terms.
Theorem 1.3.2.
A linear map $T:H_{c}\to\mathbb{C}$ is a trace on $H_{c}$ if and only if all of the following conditions hold:
@@
    $T(x^{i}R(k)e_{q})=T(y^{i}R(k)e_{q})=0$ when $i\neq 0$;
@@
    $T(x^{i}R(k)e_{q}h)=T(y^{i}R(k)e_{q}h)=0$ when $2\nmid i$;
@@
    $T(x^{2i}R(k)e_{q}h)=T(S(k)R(k+i)e_{q-i}h)$, where $S$ is the polynomial satisfying that $x^{i}y^{i}e_{q-i}=S(k)e_{q-i}$;
@@
    $T(y^{2i}R(k)e_{q}h)=(-1)^{i}T(S(k)R(k-i)e_{q+i}h)$, where $S$ is the polynomial satisfying that $y^{i}x^{i}e_{q+i}=S(k)e_{q+i}$;
@@
    $T((k+\frac{1}{2}-\alpha_{q})R(k+\frac{1}{2})e_{q})=T((k-\frac{1}{2}-\alpha_{q})R(k-\frac{1}{2})e_{q+1})$ for all $q\in\mathbb{Z}$ and $R\in\mathbb{C}[X]$;
@@
    The trace condition gives $0=T([k,x^{i}R(k)e_{q}])=-iT(x^{i}R(k)e_{q})$, and similarly $0=T([k,y^{i}R(k)e_{q}])=iT(y^{i}R(k)e_{q})$, from which the claim follows.
@@
    Conjugation by $h^{2}$ multiplies both $x$ and $y$ by 1-1- 1 while fixing $e_{q}$ and $h$, so $T(x^{i}R(k)e_{q}h)=(-1)^{i}T(x^{i}R(k)e_{q}h)$ and $T(y^{i}R(k)e_{q}h)=(-1)^{i}T(y^{i}R(k)e_{q}h)$, proving the claim.
@@
    $T(x^{2i}R(k)e_{q}h)$ $=T(x^{i}R(k)e_{q}hx^{i})$
    $=T(x^{i}R(k)e_{q}y^{i}h)$
    $=T(x^{i}y^{i}R(k+i)e_{q-i}h)$
    $=T(x^{i}y^{i}e_{q-i}R(k+i)h)$
    $=T(S(k)e_{q-i}R(k+i)h)$
    $=T(S(k)R(k+i)e_{q-i}h),$
@@
    $T(y^{2i}R(k)e_{q}h)$ $=T(y^{i}R(k)e_{q}hy^{i})$
    $=(-1)^{i}T(y^{i}R(k)e_{q}x^{i}h)$
    $=(-1)^{i}T(y^{i}x^{i}R(k-i)e_{q+i}h)$
    $=(-1)^{i}T(S(k)R(k-i)e_{q+i}h),$
@@
    The trace condition gives $0=T([k,k^{i-1}e_{q}h])=T(2k^{i}e_{q}h)$, from which the claim follows.
@@
    Using Lemma 1.2.3 and the commutativity of $e_{q}$ and $k$, we obtain
    $T\left(\left(k+\tfrac{1}{2}-\alpha_{q}\right)R\left(k+\tfrac{1}{2}\right)e_{q}\right)$ $=T\left(\left(k+\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)R\left(k+\tfrac{1}{2}\right)e_{q}\right)$
    $=\frac{1}{c_{0}}T\left(xyR\left(k+\tfrac{1}{2}\right)e_{q}\right).$
    Using the trace condition to commute $x$ and $yR(k+\frac{1}{2})e_{q}$, we find
    $T\left(\left(k+\tfrac{1}{2}-\alpha_{q}\right)R\left(k+\tfrac{1}{2}\right)e_{q}\right)$ $=\frac{1}{c_{0}}T\left(yR\left(k+\tfrac{1}{2}\right)e_{q}x\right)$
    $=\frac{1}{c_{0}}T\left(yxR\left(k-\tfrac{1}{2}\right)e_{q+1}\right)$
    $=T\left(\left(k-\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i}g^{i}}{\varepsilon^{i}-1}\right)R\left(k-\tfrac{1}{2}\right)e_{q+1}\right),$
    and using Lemma 1.2.3 and the commutativity of $e_{q}$ and $k$ once again, we obtain
    $T\left(\left(k+\tfrac{1}{2}-\alpha_{q}\right)R\left(k+\tfrac{1}{2}\right)e_{q}\right)$ $=T\left(\left(k-\tfrac{1}{2}-\alpha_{q}\right)R\left(k-\tfrac{1}{2}\right)e_{q+1}\right),$
@@
We now prove the converse. Suppose $T$ is a linear map from $H_{c}$ to $\mathbb{C}$ satisfying the listed conditions. We wish to show that for all $a$ in $H_{c}$, we have $T([x,a])=T([y,a])=T([g,a])=T([h,a])=0$. The equality $T([g,a])=T([h,a])=0$ follows from Condition (0), so it remains to verify the result for commutation by $x$ and $y$. Also, the statements that $T([y,a])=0$ for all $a$ follows from the statement that $T([x,a])=0$ for all $a$ by conjugation by $h$ and Condition (0). It therefore suffices to show that $T([x,a])=0$ when $a$ is equal to a basis element of the form $x^{i}R(k)e_{q}$, $y^{i}R(k)e_{q}$, $x^{i}R(k)e_{q}h$, or $y^{i}R(k)e_{q}h$. We now check all possible cases.
@@
  • Let $a=yR(k)e_{q}$. Then, using Lemma 1.2.3, we obtain
    $T(xa)$ $=T\left(xyR(k)e_{q}\right)$
    $=c_{0}T\left(\left(k+\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)R(k)e_{q}\right)$
    $=c_{0}T\left(\left(k+\tfrac{1}{2}-\alpha_{q}\right)R(k)e_{q}\right).$
    Using Condition (0) and Lemma 1.2.3, we obtain
    $T(xa)$ $=c_{0}T\left(\left(k-\tfrac{1}{2}-\alpha_{q}\right)R(k-1)e_{q+1}\right)$
@@
    $=T\left(yxR(k-1)e_{q+1}\right)$
    $=T\left(yR(k)e_{q}x\right)$
    $=T(ax),$
@@
    $T(xa)$ $=T\left(x^{2i}R(k)e_{q}h\right)$
    $=T\left(x^{i}y^{i}R(k+i)e_{q-i}h\right)$
    $=T\left(x^{i-1}(xy)y^{i-1}R(k+i)e_{q-i}h\right)$
    $=T\left(x^{i-1}(xy)y^{i-1}e_{q-i}R(k+i)h\right)$
    $=T\left(x^{i-1}(xy)e_{q-1}y^{i-1}R(k+i)h\right)$
    $=T\left(x^{i-1}S(k)e_{q-1}y^{i-1}R(k+i)h\right)$
    $=T\left(x^{i-1}y^{i-1}S(k+i-1)R(k+i)e_{q-i}h\right)$
    $=T\left(x^{2i-2}S(k)R(k+1)e_{q-1}h\right)$
    $=T\left(x^{2i-2}(xy)R(k+1)e_{q-1}h\right)$
    $=T\left(x^{2i-1}R(k)e_{q}yh\right)$
    $=T(x^{2i-1}R(k)e_{q}hx)$
    $=T(ax),$
@@
    $T(xa)$ $=T\left(xy^{2i+1}R(k)e_{q}h\right)$
    $=T\left(xye_{q+2i}y^{2i}R(k)h\right)$
    $=T\left(S(k)e_{q+2i}y^{2i}R(k)h\right)$
    $=T\left(y^{2i}S(k+2i)R(k)e_{q}h\right)$
    $=(-1)^{i}T\left(y^{i}x^{i}S(k+i)R(k-i)e_{q+i}h\right).$
@@
    $T(ax)$ $=T\left(y^{2i+1}R(k)e_{q}yh\right)$
    $=T\left(y^{2i+2}R(k+1)e_{q-1}h\right)$
    $=(-1)^{i+1}T\left(y^{i+1}x^{i+1}R(k-i)e_{q+i}h\right).$ (1.1)
    So by Condition (0), both $T(xa)$ and $T(ax)$ are proportional to $T(e_{q+i}h)$, which, by Lemma 1.3.1, is nonzero only if $m\mid q+i$. We may therefore assume that $m\mid q+i$, because otherwise, $T(xa)=0=T(ax)$. It follows that $n\mid 2q+2i$, so $\beta_{q+2i}=\beta_{-q}$. Therefore, by Lemma 1.2.4,
    $T(xa)$ $=(-1)^{i}T\left(y^{i}x^{i}S(k+i)R(k-i)e_{q+i}h\right)$
    $=T\left(\left(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\right)\left(c_{0}(k+i)+\beta_{q+2i}\right)R(k-i)e_{q+i}h\right)$
    $=T\left(\left(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\right)\left(c_{0}(k+i)+\beta_{-q}\right)R(k-i)e_{q+i}h\right).$
    By Condition (0), this expression depends only on its value for $k=0$. Therefore, simplifying the expression for $T(xa)$ and using Section 1.3, we obtain
    $T(xa)$ $=T\left(\left(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\right)\left(c_{0}(-k+i)+\beta_{-q}\right)R(k-i)e_{q+i}h\right)$
    $=T\left(\left(\prod\limits_{j=0}^{i}c_{0}(-k+j)+\beta_{-q-i+j}\right)R(k-i)e_{q+i}h\right)$
    $=(-1)^{i+1}T\left(y^{i+1}x^{i+1}R(k-i)e_{q+i}h\right)$
    $=T(ax),$
@@
Theorem 1.3.2 reduces traces on $H_{c}$ to traces on a subalgebra and two additional parameters, as shown by the following corollary.
Corollary 1.3.3.
A trace is uniquely determined by its values on elements of the form $R(k)e_{q}$, with $R$ a polynomial, and on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$ of $\mathbb{D}_{n}$. Also, any choice of linear function $T$ on the subalgebra of $H_{c}$ spanned by elements of the form $R(k)e_{q}$ which satisfies Conditions (0) and (0) of Theorem 1.3.2 and of values $T(g^{2i}h)$ and $T(g^{2i+1}h)$ defines a unique trace on $H_{c}$ through Conditions (0)(0) of Theorem 1.3.2.
@@
By Theorem 1.3.2, a trace is uniquely determined by its values on terms $R(k)e_{q}$ and on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$. We show that any choice of linear function on terms $R(k)e_{q}$ which satisfies Conditions (0) and (0) of Theorem 1.3.2 and of values $T(g^{2i}h)$ and $T(g^{2i+1}h)$ gives a valid trace. Indeed, values of $T$ on basis elements are determined by Conditions (0)(0) of Theorem 1.3.2, so all that is left to check is that Condition (0) of Theorem 1.3.2 does not lead to a contradiction. Using the fact that $ghg^{-1}=g^{2}h$ and $g^{i}e_{q}=\varepsilon^{-iq}e_{q}$, we see that conjugation by $g$ preserves the relations among basis elements given in Conditions (0)(0) of Theorem 1.3.2 and preserves conjugacy classes of $\mathbb{D}_{n}$. Conjugation by $h$ also preserves relationship among basis elements, so the trace is indeed well-defined. ∎
@@
By Corollary 1.3.3, it suffices to show that the dimension of the space of linear maps satisfying Conditions (0) and (0) of Theorem 1.3.2 is $\frac{n}{2}$. Let $V=\mathbb{C}[X]^{\oplus n}$, and define the linear map $\varphi\colon V\to V$ by
$\varphi(R_{0}(X),\dots,R_{n-1}(X))$
$=(R_{0}(X+\tfrac{1}{2})-R_{n-1}(X-\tfrac{1}{2}),\ \dots,\ R_{n-1}(X+\tfrac{1}{2})-R_{n-2}(X-\tfrac{1}{2})),$
where the $i$th component is equal to $R_{i}(X+\frac{1}{2})-R_{i-1}(X-\tfrac{1}{2})$. Define the ideal $I$ of $V$ by
$I=\{((X-\alpha_{0})R_{0}(X),\ \dots,\ (X-\alpha_{n-1})R_{n-1}(X)):R_{i}(X)\in\mathbb{C}[X]\}$
and the subspace $W$ of $V$ as the set of elements $(R_{0}(X),\dots,R_{n-1}(X))$ for which $R_{i}(X)=-R_{-i}(-X)$. The space of linear maps satisfying Conditions (0) and (0) of Theorem 1.3.2 is isomorphic to the vector space $(V/\langle\varphi(I),W\rangle)^{*}$, where $\langle\varphi(I),W)\rangle$ denotes the linear span of $\varphi(I)$ and $W$. We therefore wish to show that the dimension of $V/\langle\varphi(I),W\rangle$ is $\frac{n}{2}$.

### v2

@@ 4 Characterization of traces
It follows from the definition of a trace that the value of $T$ must be equal on all conjugacy classes of $\mathbb{D}_{n}$. This gives
$T(e_{q}h)={}$ $\frac{1}{n}T\left(\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{i}h\right)=\frac{1}{n}\left(\sum\limits_{i=0}^{m-1}\varepsilon^{2iq}\right)T(h)+\frac{1}{n}\left(\sum\limits_{i=0}^{m-1}\varepsilon^{(2i+1)q}\right)T(gh)$
@@
The following theorem classifies traces on $H_{c}$. In particular, conditions $(0)$ express trace values on $H_{c}$ in terms of trace values on elements of the form $R(k)e_{q}$, with $R$ a polynomial, and conditions (0) and (0) restrict possible trace values on such terms.
Theorem 4.2.
A linear map $T\colon H_{c}\to\mathbb{C}$ is a trace on $H_{c}$ if and only if all of the following conditions hold:
@@
    $T\bigl(x^{i}R(k)e_{q}\bigr)=T\bigl(y^{i}R(k)e_{q}\bigr)=0$ when $i\neq 0$;
@@
    $T\bigl(x^{i}R(k)e_{q}h\bigr)=T\bigl(y^{i}R(k)e_{q}h\bigr)=0$ when $2\nmid i$;
@@
    $T\bigl(x^{2i}R(k)e_{q}h\bigr)=T(S(k)R(k+i)e_{q-i}h)$, where $S$ is the polynomial satisfying that $x^{i}y^{i}e_{q-i}\allowbreak=S(k)e_{q-i}$;
@@
    $T\bigl(y^{2i}R(k)e_{q}h\bigr)=(-1)^{i}T(S(k)R(k-i)e_{q+i}h)$, where $S$ is the polynomial satisfying that $y^{i}x^{i}e_{q+i}=S(k)e_{q+i}$;
@@
    $T\bigl(\bigl(k+\frac{1}{2}-\alpha_{q}\bigr)R\bigl(k+\frac{1}{2}\bigr)e_{q}\bigr)=T\bigl(\bigl(k-\frac{1}{2}-\alpha_{q}\bigr)R\bigl(k-\frac{1}{2}\bigr)e_{q+1}\bigr)$ for all $q\in\mathbb{Z}$ and $R\in\mathbb{C}[X]$;
@@
    The trace condition gives $0=T\bigl(\big[k,x^{i}R(k)e_{q}\big]\bigr)=-iT\bigl(x^{i}R(k)e_{q}\bigr)$, and similarly $0=T\bigl(\big[k,y^{i}R(k)e_{q}\big]\bigr)=iT\bigl(y^{i}R(k)e_{q}\bigr)$, from which the claim follows.
@@
    Conjugation by $h^{2}$ multiplies both $x$ and $y$ by 1-1 while fixing $e_{q}$ and $h$, so
    $T\bigl(x^{i}R(k)e_{q}h\bigr)=(-1)^{i}T\bigl(x^{i}R(k)e_{q}h\bigr)\qquad\text{and}$
    $T\bigl(y^{i}R(k)e_{q}h\bigr)=(-1)^{i}T\bigl(y^{i}R(k)e_{q}h\bigr),$
    proving the claim.
@@
    $T\bigl(x^{2i}R(k)e_{q}h\bigr)={}$ $T\bigl(x^{i}R(k)e_{q}hx^{i}\bigr)=T\bigl(x^{i}R(k)e_{q}y^{i}h\bigr)$
    $={}$ $T\bigl(x^{i}y^{i}R(k+i)e_{q-i}h\bigr)=T\bigl(x^{i}y^{i}e_{q-i}R(k+i)h\bigr)$
    $={}$ $T(S(k)e_{q-i}R(k+i)h)=T(S(k)R(k+i)e_{q-i}h),$
@@
    $T\bigl(y^{2i}R(k)e_{q}h\bigr)={}$ $T\bigl(y^{i}R(k)e_{q}hy^{i}\bigr)=(-1)^{i}T\bigl(y^{i}R(k)e_{q}x^{i}h\bigr)$
    $={}$ $(-1)^{i}T\bigl(y^{i}x^{i}R(k-i)e_{q+i}h\bigr)=(-1)^{i}T(S(k)R(k-i)e_{q+i}h),$
@@
    The trace condition gives $0=T\bigl(\big[k,k^{i-1}e_{q}h\big]\bigr)=T\bigl(2k^{i}e_{q}h\bigr)$, from which the claim follows.
@@
    Using Lemma 3.3 and the commutativity of $e_{q}$ and $k$, we obtain
    $T\bigl(\bigl(k+\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k+\tfrac{1}{2}\bigr)e_{q}\bigr)={}$ $T\left(\left(k+\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)R\bigl(k+\tfrac{1}{2}\bigr)e_{q}\right)$
    $={}$ $\frac{1}{c_{0}}T\bigl(xyR\bigl(k+\tfrac{1}{2}\bigr)e_{q}\bigr).$
    Using the trace condition to commute $x$ and $yR\bigl(k+\frac{1}{2}\bigr)e_{q}$, we find
    $T\bigl(\bigl(k+\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k+\tfrac{1}{2}\bigr)e_{q}\bigr)={}$ $\frac{1}{c_{0}}T\bigl(yR\bigl(k+\tfrac{1}{2}\bigr)e_{q}x\bigr)=\frac{1}{c_{0}}T\bigl(yxR\bigl(k-\tfrac{1}{2}\bigr)e_{q+1}\bigr)$
    $={}$ $T\left(\left(k-\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i}g^{i}}{\varepsilon^{i}-1}\right)R\bigl(k-\tfrac{1}{2}\bigr)e_{q+1}\right),$
    and using Lemma 3.3 and the commutativity of $e_{q}$ and $k$ once again, we obtain
    $T\bigl(\bigl(k+\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k+\tfrac{1}{2}\bigr)e_{q}\bigr)={}$ $T\bigl(\bigl(k-\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k-\tfrac{1}{2}\bigr)e_{q+1}\bigr),$
@@
We now prove the converse. Suppose $T$ is a linear map from $H_{c}$ to $\mathbb{C}$ satisfying the listed conditions. We wish to show that for all $a$ in $H_{c}$, we have
$T([x,a])=T([y,a])=T([g,a])=T([h,a])=0.$
The equality $T([g,a])=T([h,a])=0$ follows from condition (0), so it remains to verify the result for commutation by $x$ and $y$. Also, the statements that $T([y,a])=0$ for all $a$ follows from the statement that $T([x,a])=0$ for all $a$ by conjugation by $h$ and condition (0). It therefore suffices to show that $T([x,a])=0$ when $a$ is equal to a basis element of the form $x^{i}R(k)e_{q}$, $y^{i}R(k)e_{q}$, $x^{i}R(k)e_{q}h$, or $y^{i}R(k)e_{q}h$. We now check all possible cases.
@@
  • Let $a=yR(k)e_{q}$. Then, using Lemma 3.3, we obtain
    $T(xa)={}$ $T\left(xyR(k)e_{q}\right)=c_{0}T\left(\left(k+\frac{1}{2}-\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)R(k)e_{q}\right)$
    $={}$ $c_{0}T\bigl(\bigl(k+\tfrac{1}{2}-\alpha_{q}\bigr)R(k)e_{q}\bigr).$
    Using condition (0) and Lemma 3.3, we obtain
    $T(xa)={}$ $c_{0}T\bigl(\bigl(k-\tfrac{1}{2}-\alpha_{q}\bigr)R(k-1)e_{q+1}\bigr)$
@@
    $={}$ $T\left(yxR(k-1)e_{q+1}\right)=T\left(yR(k)e_{q}x\right)=T(ax),$
@@
    $T(xa)={}$ $T\bigl(x^{2i}R(k)e_{q}h\bigr)=T\bigl(x^{i}y^{i}R(k+i)e_{q-i}h\bigr)$
    $={}$ $T\bigl(x^{i-1}(xy)y^{i-1}R(k+i)e_{q-i}h\bigr)=T\bigl(x^{i-1}(xy)y^{i-1}e_{q-i}R(k+i)h\bigr)$
    $={}$ $T\bigl(x^{i-1}(xy)e_{q-1}y^{i-1}R(k+i)h\bigr)=T\bigl(x^{i-1}S(k)e_{q-1}y^{i-1}R(k+i)h\bigr)$
    $={}$ $T\bigl(x^{i-1}y^{i-1}S(k+i-1)R(k+i)e_{q-i}h\bigr)=T\bigl(x^{2i-2}S(k)R(k+1)e_{q-1}h\bigr)$
    $={}$ $T\bigl(x^{2i-2}(xy)R(k+1)e_{q-1}h\bigr)=T\bigl(x^{2i-1}R(k)e_{q}yh\bigr)$
    $={}$ $T\bigl(x^{2i-1}R(k)e_{q}hx\bigr)=T(ax),$
@@
    $T(xa)={}$ $T\bigl(xy^{2i+1}R(k)e_{q}h\bigr)=T\bigl(xye_{q+2i}y^{2i}R(k)h\bigr)$
    $={}$ $T\bigl(S(k)e_{q+2i}y^{2i}R(k)h\bigr)=T\bigl(y^{2i}S(k+2i)R(k)e_{q}h\bigr)$
    $={}$ $(-1)^{i}T\bigl(y^{i}x^{i}S(k+i)R(k-i)e_{q+i}h\bigr).$
@@
    $T(ax)={}$ $T\bigl(y^{2i+1}R(k)e_{q}yh\bigr)=T\bigl(y^{2i+2}R(k+1)e_{q-1}h\bigr)$
    $={}$ $(-1)^{i+1}T\bigl(y^{i+1}x^{i+1}R(k-i)e_{q+i}h\bigr).$ (4.1)
    So by condition (0), both $T(xa)$ and $T(ax)$ are proportional to $T(e_{q+i}h)$, which, by Lemma 4.1, is nonzero only if $m\mid q+i$. We may therefore assume that $m\mid q+i$, because otherwise, $T(xa)=0=T(ax)$. It follows that $n\mid 2q+2i$, so $\beta_{q+2i}=\beta_{-q}$. Therefore, by Lemma 3.4,
    $T(xa)={}$ $(-1)^{i}T\bigl(y^{i}x^{i}S(k+i)R(k-i)e_{q+i}h\bigr)$
    $\ ={}$ $T\Bigg(\Bigg(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\Bigg)(c_{0}(k+i)+\beta_{q+2i})R(k-i)e_{q+i}h\Bigg)$
    $={}$ $T\Bigg(\Bigg(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\Bigg)(c_{0}(k+i)+\beta_{-q})R(k-i)e_{q+i}h\Bigg).$
    By condition (0), this expression depends only on its value for $k=0$. Therefore, simplifying the expression for $T(xa)$ and using (4.1), we obtain
    $T(xa)={}$ $T\Bigg(\Bigg(\prod\limits_{j=0}^{i-1}c_{0}(-k+j)+\beta_{-q-i+j}\Bigg)(c_{0}(-k+i)+\beta_{-q})R(k-i)e_{q+i}h\Bigg)$
    $={}$ $T\Bigg(\Bigg(\prod\limits_{j=0}^{i}c_{0}(-k+j)+\beta_{-q-i+j}\Bigg)R(k-i)e_{q+i}h\Bigg)$
    $={}$ $(-1)^{i+1}T\bigl(y^{i+1}x^{i+1}R(k-i)e_{q+i}h\bigr)=T(ax),$
@@
Theorem 4.2 reduces traces on $H_{c}$ to traces on a subalgebra and two additional parameters, as shown by the following corollary.
Corollary 4.3.
A trace is uniquely determined by its values on elements of the form $R(k)e_{q}$, with $R$ a polynomial, and on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$ of $\mathbb{D}_{n}$. Also, any choice of linear function $T$ on the subalgebra of $H_{c}$ spanned by elements of the form $R(k)e_{q}$ which satisfies conditions (0) and (0) of Theorem 4.2 and of values $T\bigl(g^{2i}h\bigr)$ and $T\bigl(g^{2i+1}h\bigr)$ defines a unique trace on $H_{c}$ through conditions $(0)$ of Theorem 4.2.
@@
By Theorem 4.2, a trace is uniquely determined by its values on terms $R(k)e_{q}$ and on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$. We show that any choice of linear function on terms $R(k)e_{q}$ which satisfies conditions (0) and (0) of Theorem 4.2 and of values $T\bigl(g^{2i}h\bigr)$ and $T\bigl(g^{2i+1}h\bigr)$ gives a valid trace. Indeed, values of $T$ on basis elements are determined by conditions $(0)$ of Theorem 4.2, so all that is left to check is that condition (0) of Theorem 4.2 does not lead to a contradiction. Using the fact that $ghg^{-1}=g^{2}h$ and $g^{i}e_{q}=\varepsilon^{-iq}e_{q}$, we see that conjugation by $g$ preserves the relations among basis elements given in conditions $(0)$ of Theorem 4.2 and preserves conjugacy classes of $\mathbb{D}_{n}$. Conjugation by $h$ also preserves relationship among basis elements, so the trace is indeed well-defined. ∎
@@
By Corollary 4.3, it suffices to show that the dimension of the space of linear maps satisfying conditions (0) and (0) of Theorem 4.2 is $\frac{n}{2}$. Let $V=\mathbb{C}[X]^{\oplus n}$, and define the linear map $\varphi\colon V\to V$ by
$\begin{split}&\varphi(R_{0}(X),\dots,R_{n-1}(X))\\ &\qquad=\bigl(R_{0}\bigl(X+\tfrac{1}{2}\bigr)-R_{n-1}\bigl(X-\tfrac{1}{2}\bigr),\dots,R_{n-1}\bigl(X+\tfrac{1}{2}\bigr)-R_{n-2}\bigl(X-\tfrac{1}{2}\bigr)\bigr),\end{split}$
where the $i$th component is equal to $R_{i}\bigl(X+\frac{1}{2}\bigr)-R_{i-1}\bigl(X-\tfrac{1}{2}\bigr)$. Define the ideal $I$ of $V$ by
$I=\{((X-\alpha_{0})R_{0}(X),\dots,(X-\alpha_{n-1})R_{n-1}(X))\mid R_{i}(X)\in\mathbb{C}[X]\}$
and the subspace $W$ of $V$ as the set of elements $(R_{0}(X),\dots,R_{n-1}(X))$ for which $R_{i}(X)=-R_{-i}(-X)$. The space of linear maps satisfying conditions (0) and (0) of Theorem 4.2 is isomorphic to the vector space $(V/\langle\varphi(I),W\rangle)^{*}$, where $\langle\varphi(I),W)\rangle$ denotes the linear span of $\varphi(I)$ and $W$. We therefore wish to show that the dimension of $V/\langle\varphi(I),W\rangle$ is $\frac{n}{2}$.